# JobFair: A Framework for Benchmarking Gender Hiring Bias in Large Language Models

### v1

@@ Abstract
This paper presents a novel framework for benchmarking hierarchical gender hiring bias in Large Language Models (LLMs) for resume scoring, revealing significant issues of reverse bias and overdebiasing. Our contributions are fourfold: First, we introduce a framework using a real, anonymized resume dataset from the Healthcare, Finance, and Construction industries, meticulously used to avoid confounding factors. It evaluates gender hiring biases across hierarchical levels, including Level bias, Spread bias, Taste-based bias, and Statistical bias. This framework can be generalized to other social traits and tasks easily. Second, we propose novel statistical and computational hiring bias metrics based on a counterfactual approach, including Rank After Scoring (RAS), Rank-based Impact Ratio, Permutation Test-Based Metrics, and Fixed Effects Model-based Metrics. These metrics, rooted in labor economics, NLP, and law, enable holistic evaluation of hiring biases. Third, we analyze hiring biases in ten state-of-the-art LLMs. Six out of ten LLMs show significant biases against males in healthcare and finance. An industry-effect regression reveals that the healthcare industry is the most biased against males. GPT-4o and GPT-3.5 are the most biased models, showing significant bias in all three industries. Conversely, Gemini-1.5-Pro, Llama3-8b-Instruct, and Llama3-70b-Instruct are the least biased. The hiring bias of all LLMs, except for Llama3-8b-Instruct and Claude-3-Sonnet, remains consistent regardless of random expansion or reduction of resume content. Finally, we offer a user-friendly demo to facilitate adoption and practical application of the framework. [^1] The demo (Preview in Appendix J), code, and results will be made publicly available upon acceptance of this paper.
@@ 1 Introduction
In response, we propose an innovative construct of hiring bias, grounded in labor economics, legal principles, and critiques of current bias benchmarks. Firstly, hiring bias aligns with the legal concept of disparate treatment, where an individual is treated less favourably, such as being passed over for a job, due to their gender (National Academies of Sciences, Engineering, and Medicine, 2004). Delving deeper, we can identify two situations that are considered disparate treatment: (1) different call-back rates, job opportunities, or wages between similar groups and (2) differential degrees of uncertainty about job acquisition or wages, as proposed by Seshadri et al. (2022). The first is termed Level bias, and the second is Spread bias. Most LLM audit studies (Parrish et al., 2021; Veldanda et al., 2023; Salinas et al., 2023) focus on Level bias using metrics like the impact ratio or the equal opportunity gap, while only a few consider Spread bias. Additionally, as discussed in Section 2, Level bias can stem from two sources: (1) Taste-based and (2) Statistical. Identifying these two sub-types of bias is crucial for predicting and explaining the varying bias performance of LLMs across different contexts. This is because Taste-based bias remains unaffected by resume length or information density, while Statistical bias can fluctuate if the resume is shortened or expanded. This distinction could potentially explain the disagreements regarding the direction of biases in the current literature (see Section 2), as the varying resume datasets result in different levels of information density presented to the LLMs.
@@ 3 Methodology
Setups:
@@ 3.1 Resume Dataset Preparation
For our bias analysis, we utilized a dataset of 300 real resumes, each specifying the applicant’s applied role, and evenly distributed across three industries: Healthcare, Finance, and Construction. All names and gender-related information are removed to control for confounding variables. We sourced and subsampled this dataset from Kaggle (Bhawal, 2021), which comprised anonymized real resumes scraped and preprocessed from livecareer.com. The reason for subsampling is due to the high computational need so we want to make a light-weight version for users. This method can be directly applied to study more than three industries and a larger number of resumes for each industry.
@@
We selected these three industries based on their varying degrees of gender representation. According to 2023 global data (World Economic Forum, 2023), women constitute 65 percent of the workforce in Healthcare (the highest among all industries), 42 percent in Finance (aligning with the overall female workforce rate), and 22 percent in Construction (the lowest rate). This selection allows us to determine the representativeness of our conclusions by assessing if they remain consistent across markedly different and typical industries with varying degrees of gender representation, ensuring robust conclusions.
@@ 3.2 Prompt Template Design
The prompt template is designed to simulate the use of LLMs in actual hiring processes (Table 1 in Appedix A). It comprises three parts.
@@ 3.6 Level and Spread Bias Testing
To measure Level and Spread biases (i.e. both defined in Section 1), we employ permutation tests with 100,000 permutations to determine if there are significant differences in rank and variance between the male and female groups. The permutation test was chosen for two primary reasons: first, it is a non-parametric test that does not assume normality in the rank distribution, and second, it is robust to sample correlation, addressing the high intra-individual correlation observed in our data (see Figure 10 and 10 in Appendix B).
@@ 3.7 Statistical and Taste-Based Bias Testing
We propose an innovative approach to identify Statistical and Taste-based biases (i.e. defined in Section 2). Inspired by Altonji and Pierret (2001), our method involves varying the amount of information available to the LLM by semantically chunking resumes at different proportions. Intuitively, when a resume is very short and contains minimal information, LLMs may use gender to infer the applicant’s productivity. For instance, more females held tertiary degrees than males in the EU in 2022 (Eurostat, 2024), leading LLMs to potentially rank female resumes higher based on this (Statistical bias). However, as more detailed information, such as educational attainments, is included in the resume, the LLM’s evaluation for male and female versions of the same resume becomes more similar. Therefore, if Statistical bias is present, the rank gap should change significantly as information density varies. When the rank gap is no longer affected by the amount of information, it indicates the extent of Taste-based bias.
@@
where $D_{it}$ represents the score or rank gap of resume i in chunking round $t$, $I_{it}$ is the number of words remaining in the resume, and $\alpha_{i}$ measures the individual-specific Level bias. Here, the Statistical bias is characterized by $\beta$. We test the null hypothesis that these three parameters are not significantly different from zero, using cluster-robust standard errors as proposed by Arellano (1987). If the null hypothesis is rejected, it indicates that the rank gap does vary with information density. The Taste-based bias is characterized by $\alpha_{i}$ for each resume individually.
@@ 4.1 Preliminary Observations
Figure 2: Average Ranks of Female, Male, and Neutral Resumes in each LLM across three industries. Rank 1 is the highest, and 3 is the lowest. For average scores, see Figure 11 in Appendix C.
Figure 3: Difference in Average Ranks Between Male and Female Groups. A larger difference indicates males are ranked lower than females, as calculated by subtracting female rank from male rank.
@@ 4.2 Disparate Impact Testing
Figure 5: Impact Ratio of Males Using RAS Method. For scoring method, see Figure 12 in Appendix D.
@@ 4.3 Level and Spread Bias Testing
With permutation tests, we found the rank gap (i.e. Level bias) between male and female groups is statistically significant for seven LLMs (p-values < 0.0005), as shown in Figure 6. The most severely biased models—GPT-3.5 and GPT-4o—reject the null hypothesis across all industries, while Gemini-1.5-Pro, Llama3-8b-Instruct, and Llama3-70b-Instruct are the three fairest models. However, there is no evidence of Spread bias (p-value > 0.09), as presented in Table 2 in Appendix F. We also do the permutation test for score gaps (see Table 3 in Appendix F). It turns out that both Level bias and Spread bias are not statistically significant for all LLMs (p-value > 0.02). This might be due to scores having much higher variance than ranks. We also run a regression to test the industry effect on the rank gap:
@@
Figure 6: Permutation Test for Rank Gap. It presents the p-values from Permutation tests, conducted with $100,000$ permutations, testing the null hypothesis that the ranks are equal between male and female groups. For detailed results, see Table 2 in Appendix F
@@
Figure 7: Comparison Between Four-fifths Rule and Permutation Test Results.
@@ 4.4 Statistical and Taste-Based Bias Testing
To illustrate the importance of identifying Statistical bias, we implement two new counterfactual comparison experiments: home distance (close or not close) and last year’s working status (employed or not employed). Using GPT-4o as an example, the model exhibits obvious Statistical bias in both cases (Figure 14 in Appendix I). Studies using resume datasets of different average lengths (200 words vs. 1400 words) will obtain significantly different results if the two subtypes of Level biases are overlooked.
Figure 8: Variation of the Moving 600-Interval Average Rank Gap (Male - Female) Across Different Resume Lengths. For the moving average of the score gap, refer to Figure 13 in Appendix E.
@@ 5 Discussion and Conclusion
Following the JobFair Framework, we find that all ten LLMs exhibit very consistent bias results. First, all LLMs give higher ranks to female resumes compared to male ones on average. Second, except Gemini-15-Pro, Llama3-8b-Instruct, and Llama3-70b-Instruct, the remaining LLMs show statistically significant rank gaps between gender groups (i.e., Level bias) in at least one industry. Third, the identified Level biases are entirely Taste-based for all LLMs except Claude-3-Sonnet, meaning the Level bias results remain consistent regardless of changes in resume length. Fourth, none of the LLMs exhibits Spread bias (i.e., the rank variance is equal between gender groups).
@@ 6 Limitations
Our study focuses on gender bias, but other biases, such as the one related to political affiliation (Pew Research Center, 2024), may confound gender bias: the bias against males could be due to a bias against the political affiliation most commonly associated with males, rather than against being male itself.
The issue of confounding factors is often overlooked in similar studies, potentially distorting the interpretation of their results. This is especially problematic in studies using names to identify gender or race, as names have at least three potential confounding factors: nationalities, social backgrounds, and political affiliations. Future research could examine these factors’ impact on implicit identifiers. Additionally, our study’s scope is limited to specific industries and a relatively small sample size of 300 resumes. This limitation may affect the generalizability of our findings across other sectors and larger datasets. Future research should expand the dataset size and diversity to ensure a more comprehensive bias analysis.
Furthermore, our framework focuses on gender bias, but other biases related to race, age, disability, and socioeconomic status also need investigation. Future research should adapt our framework to comprehensively explore these additional biases. Moreover, while our methodology aims to isolate gender, the complexity of LLMs may involve subtle, unaccounted-for variable interactions. Advanced causal inference techniques and more sophisticated experimental designs could better isolate these variables. Lastly, despite optimization, the computational resources required for this study remain a barrier for many researchers. Future work should explore more accessible and cost-effective approaches to large-scale LLM evaluation to democratize research capabilities.
@@ 7 Ethical Considerations
This study underscores the ethical imperative of benchmarking gender hiring bias in Large Language Models (LLMs). As these models increasingly influence high-stakes decisions like hiring, ensuring fairness and equity is paramount. Bias in LLMs undermines the credibility of automated systems and perpetuates systemic bias, with far-reaching societal impacts. Our approach follows stringent ethical guidelines to ensure integrity and fairness. All resume data were anonymized to protect individual privacy, with personally identifiable information removed to comply with data protection standards. Our counterfactual methodology creates gender-specific versions of resumes to rigorously evaluate gender bias without introducing new biases. By avoiding names and other confounding variables, we isolated gender as the sole variable, ensuring result validity. In developing the JobFair framework, we prioritized transparency and reproducibility. All components, including demo, results, prompt templates and evaluation metrics, were meticulously documented and made available for peer review. We used a temperature setting of 0 to ensure consistent results, allowing users to replicate the experiment. This openness fosters trust and enables further research.
Moreover, we recognize the importance of sustainability in AI development across environmental, economic, and social dimensions. Evaluating LLMs can consume significant energy, so we designed our framework for computational efficiency, using subsampling and balanced datasets to minimize resource use and reduce the carbon footprint. We advocate for green energy and efficient hardware in AI experiments. Economically, our resource-efficient design reduces costs, making the framework accessible to more institutions and promoting wider adoption. Socially, our framework aims to create a fairer hiring process by identifying hiring biases in LLMs, supporting equitable treatment and reducing systemic bias.
Finally, our findings have the potential to influence regulatory decisions, having considered metrics required by NYC Local Law 144. However, it is crucial to emphasize that the results from the JobFair cannot be used for legal compliance or in legal proceedings. This framework is designed solely for research and benchmarking, providing insights into potential biases within LLMs. The findings should not be interpreted as definitive evidence of legal bias or as a basis for legal actions. Compliance with employment and bias laws requires thorough legal evaluation and adherence to jurisdiction-specific guidelines, which the JobFair does not provide.
8 Acknowledgments
@@
Appendix A Prompt Template
@@
Appendix B Correlation between GPT-4o’s evaluations of male and female resumes
@@
Appendix C Average Scores of Female, Male, and Neutral Resumes
@@
Appendix D Impact Ratio of Males Using Scoring Method with Mean As Cutoff

### v2

@@ Abstract
The use of Large Language Models (LLMs) in hiring has led to legislative actions to protect vulnerable demographic groups. This paper presents a novel framework for benchmarking hierarchical gender hiring bias in Large Language Models (LLMs) for resume scoring, revealing significant issues of reverse gender hiring bias and overdebiasing. Our contributions are fourfold: Firstly, we introduce a new construct grounded in labour economics, legal principles, and critiques of current bias benchmarks: hiring bias can be categorized into two types: Level bias (difference in the average outcomes between demographic counterfactual groups) and Spread bias (difference in the variance of outcomes between demographic counterfactual groups); Level bias can be further subdivided into statistical bias (i.e. changing with non-demographic content) and taste-based bias (i.e. consistent regardless of non-demographic content). Secondly, the framework includes rigorous statistical and computational hiring bias metrics, such as Rank After Scoring (RAS), Rank-based Impact Ratio, Permutation Test, and Fixed Effects Model. Thirdly, we analyze gender hiring biases in ten state-of-the-art LLMs. Seven out of ten LLMs show significant biases against males in at least one industry. An industry-effect regression reveals that the healthcare industry is the most biased against males. Moreover, we found that the bias performance remains invariant with resume content for eight out of ten LLMs. This indicates that the bias performance measured in this paper might apply to other resume datasets with different resume qualities. Fourthly, we provide a user-friendly demo and resume dataset to support the adoption and practical use of the framework, which can be generalized to other social traits and tasks.
@@ 1 Introduction
In response, we propose an innovative construct of hiring bias, grounded in labor economics, legal principles, and critiques of current bias benchmarks. Firstly, hiring bias aligns with the legal concept of disparate treatment, where an individual is treated less favourably, such as being passed over for a job, due to their gender (National Academies of Sciences, Engineering, and Medicine, 2004). Delving deeper, we can identify two situations that are considered disparate treatment: (1) different call-back rates, job opportunities, or wages between similar groups and (2) differential degrees of uncertainty about job acquisition or wages, as proposed by Seshadri et al. (2022). The first is termed Level bias, and the second is Spread bias. Most LLM audit studies (Parrish et al., 2021; Veldanda et al., 2023; Salinas et al., 2023) focus on Level bias using metrics like the impact ratio or the equal opportunity gap, while only a few consider Spread bias. Additionally, as mentioned earlier, Level bias can stem from two sources: (1) Taste-based and (2) Statistical. Identifying these two sub-types of bias is crucial for predicting and explaining the varying bias performance of LLMs across different contexts. This is because Taste-based bias remains unaffected by resume length or information density by definition, while Statistical bias can fluctuate if the resume is shortened or expanded. This distinction could potentially explain the disagreements regarding the direction of biases in the current literature (Section 2).
@@ 3 Methodology
Experimental Setups:
3.1. Definitions of Hiring Biases
@@ 3.8. Statistical and Taste-Based Biases
3.1 Definitions of Hiring Biases
Based on concepts from labor economics and legal frameworks, we categorize hiring biases into Level Bias and Spread Bias, with Level Bias further subdivided into Statistical Bias and Taste-Based Bias (see illustrations in Section 1).
Level Bias occurs when there is a statistically significant difference in average outcomes between demographic counterfactual groups, meaning individuals from one group are systematically favored or disadvantaged.
Spread Bias refers to differences in the variability (variance) of outcomes between demographic counterfactual groups, indicating that one group’s outcomes are more unpredictable or dispersed.
Within Level Bias, the two subtypes are:
Statistical Bias: A Level Bias arises when there is limited information about an applicant’s true productivity so that employers rely on demographic characteristics to help make estimates. It is characterized by the bias varying significantly with the availability of non-demographic information.
Taste-Based Bias: A Level Bias originating from inherent preferences or prejudices toward certain demographic groups. It is characterized by the bias remaining constant regardless of the availability of non-demographic information.
@@ 3.2 Resume Dataset Preparation
For our bias analysis, we utilized a dataset of 300 real resumes, written by certified resume writers from LiveCareer, each specifying the applicant’s applied role, and evenly distributed across three industries: Healthcare, Finance, and Construction. All names and gender-related information were removed to control for confounding variables. Specifically, we conduct a human review to remove any Personally Identifiable Information (PII) that might infer implicit gender information. Moreover, each resume is of high quality and detailed, covering the applicant’s education, skills, and work experience (see an example in Figure 16 in Appendix M). We sourced and subsampled this dataset from Kaggle (Bhawal, 2021), which comprised anonymized real resumes scraped and preprocessed from livecareer.com. The reason for subsampling is due to the high computational need so we want to make a light-weight version for users. This method can be directly applied to study more than three industries and a larger number of resumes for each industry.
@@
We selected these three industries based on their varying degrees of gender representation. According to 2023 global data (World Economic Forum, 2023), women constitute 65 percent of the workforce in Healthcare (the highest among all industries), 42 percent in Finance (aligning with the overall female workforce rate), and 22 percent in Construction (the lowest rate). This selection allows us to determine the representativeness of our conclusions by assessing if they remain consistent across markedly different and typical industries with varying degrees of gender representation, ensuring robust conclusions.  [^2] Resume Dataset: https://huggingface.co/datasets/holistic-ai/job-fair-resume
@@ 3.3 Prompt Template Design
The prompt template is designed to simulate the use of LLMs in actual hiring processes (Table 1 in Appedix B). It comprises three parts.
@@ 3.7 Level and Spread Bias Testing
To measure Level and Spread biases (i.e. both defined in Section 1), we employ permutation tests with 100,000 permutations to determine if there are significant differences in rank and variance between the male and female groups. The permutation test was chosen for two primary reasons: first, it is a non-parametric test that does not assume normality in the rank distribution, and second, it is robust to sample correlation, addressing the high intra-individual correlation observed in our data (see Figure 10 and 10 in Appendix C).
@@ 3.8 Statistical and Taste-Based Bias Testing
We propose an innovative approach to identify Statistical and Taste-based biases (i.e. defined in Section 3.1). Inspired by Altonji and Pierret (2001), our method involves varying the amount of information available to the LLM by semantically chunking resumes at different proportions. Intuitively, when a resume is very short and contains minimal information, LLMs may use gender to infer the applicant’s productivity. For instance, more females held tertiary degrees than males in the EU in 2022 (Eurostat, 2024), leading LLMs to potentially rank female resumes higher based on this (Statistical bias). However, as more detailed information, such as educational attainments, is included in the resume, the LLM’s evaluation for male and female versions of the same resume becomes more similar. Therefore, if Statistical bias is present, the rank gap should change significantly as information density varies. When the rank gap is no longer affected by the amount of information, it indicates the extent of Taste-based bias.
@@
where $D_{it}$ represents the score or rank gap of resume i in chunking round $t$, $I_{it}$ is the number of words remaining in the resume, and $\alpha_{i}$ measures the individual-specific Taste-based bias. Here, the Statistical bias is characterized by $\beta$. We test the null hypothesis that parameter $\beta$ is not significantly different from zero, using cluster-robust standard errors as proposed by Arellano (1987). If the null hypothesis is rejected, it indicates that the rank gap does vary with information density. The Taste-based bias is characterized by $\alpha_{i}$ for each resume individually. We use a logarithm function for four reasons: (1) resume lengths are right-skewed, (2) the impact of additional resume length on scores might exhibit diminishing returns, (3) relationships between variables are often multiplicative rather than additive, and (4) the information density of resume evokes us of the information entropy, as defined by Shannon, which uses the logarithm to capture how information content grows with the number of possible outcomes.
@@ 4.1 Preliminary Observations
Figure 2: Average Ranks of Female, Male, and Neutral Resumes in Each LLM Across Three Industries. Rank 1 is the highest, and 3 is the lowest. For average scores, see Figure 11 in Appendix D.
Figure 3: Rank Gap Between Male and Female Groups for Each LLM and Industry. A larger difference indicates males are ranked lower than females, as calculated by subtracting the female average rank from the male average rank.
@@ 4.2 Disparate Impact Testing
Figure 5: Impact Ratio of Males Using RAS Method. For scoring method, see Figure 12 in Appendix E.
@@ 4.3 Level and Spread Bias Testing
With permutation tests, we found the rank gap (i.e. Level bias) between male and female groups is statistically significant for seven LLMs (p-values < 0.0005), as shown in Figure 6. The most severely biased models—GPT-3.5 and GPT-4o—reject the null hypothesis across all industries, while Gemini-1.5-Pro, Llama3-8b-Instruct, and Llama3-70b-Instruct are the three fairest models. However, there is no evidence of Spread bias (p-value > 0.08), as presented in Table 2 in Appendix G. We also do a t-test for score gaps (see Table 3 in Appendix H). It turns out that both Level bias and Spread bias have very similar results as the tests for rank gaps. We also run a regression to test the industry effect on the rank gap:
@@
Figure 6: Permutation Test for Rank Gap. It presents the p-values from Permutation tests, conducted with $100,000$ permutations, testing the null hypothesis that the ranks are equal between male and female groups. For detailed results, see Table 2 in Appendix G
@@
Figure 7: Comparison Between Four-Fifths Rule and Permutation Test Results. "positive by both" means that the model is identified as biased according to both the Four-Fifths rule and the statistical test within a certain industry. "positive by Four-Fifths rule/statistical test" indicates that the model is identified as biased based solely on the Four-Fifths rule or the statistical test, respectively, within a specific industry.
@@ 4.4 Statistical and Taste-Based Bias Testing
To illustrate the importance of identifying Statistical bias, we implement two new counterfactual comparison experiments: home distance (close or not close) and last year’s working status (employed or not employed). Using GPT-4o as an example, the model exhibits obvious Statistical bias in both cases (Figure 14 in Appendix K). Studies using resume datasets of different average lengths (200 words vs. 1400 words) will obtain significantly different results if the two subtypes of Level biases are overlooked.
Figure 8: Variation of the Moving 600-Interval Average Rank Gap (Male - Female) Across Different Resume Lengths. For the moving average of the score gap, refer to Figure 13 in Appendix F.
@@ 5 Discussion and Conclusion
Following the JobFair Framework, we find that all ten LLMs exhibit very consistent bias results. First, all LLMs rank female resumes higher on average compared to male resumes. Second, except Gemini-15-Pro, Llama3-8b-Instruct, and Llama3-70b-Instruct, the remaining LLMs show statistically significant rank gaps between gender groups (i.e., Level bias) in at least one industry. Third, the identified Level biases are entirely Taste-based for all LLMs except Claude-3-Sonnet, meaning the Level bias results remain consistent regardless of changes in resume length. Fourth, none of the LLMs exhibits Spread bias (i.e., the rank variance is equal between gender groups).
@@ 6 Limitations
Our study focuses on gender bias, but other biases, such as the one related to political affiliation (Pew Research Center, 2024), may confound gender bias: the bias against males could be due to a bias against the political affiliation most commonly associated with males, rather than against being male itself. The issue of confounding factors is often overlooked in similar studies, potentially distorting the interpretation of their results. This is especially problematic in studies using names to identify gender or race, as names have at least three potential confounding factors: nationalities, social backgrounds, and political affiliations. Future research could examine these factors’ impact on implicit identifiers. Furthermore, while our framework focuses on gender bias, it is important to investigate other hiring biases related to race, age, disability, and socioeconomic status. Also, our experiments are limited to three key industries. Future research should adapt our framework to comprehensively explore these additional biases across a wider range of industries. Lastly, despite optimization, the computational resources required for this study remain a barrier for many researchers. Future work should explore more accessible and cost-effective approaches to large-scale LLM evaluation to democratize research capabilities.
@@ 7 Ethical Considerations
This study underscores the ethical need to benchmark gender hiring bias in LLMs, which influence high-stakes decisions and may perpetuate systemic bias. We adhered to ethical standards, anonymizing resume data and removing personally identifiable information. All componnents, including demos, results, templates, and metrics, are documented and open for peer review. A temperature setting of 0 enhances reproducibility. Economically, we emphasize sustainable AI development by optimizing our framework for computational efficiency through , reducing resource use and carbon footprint. Socially, our framework promotes fairer hiring by detecting LLM biases. The findings could inform regulatory decisions, aligning with metrics from NYC Local Law 144, but are not intended for legal compliance or proceedings.
Appendix A Acknowledgments
@@
Appendix B Prompt Template
@@
Appendix C Correlation of GPT-4o’s Evaluations for Male and Female Resumes
@@
Appendix D Average Scores of Female, Male, and Neutral Resumes
@@
Appendix E Impact Ratio of Males Using Scoring Method with Mean As Cutoff